OpenStudy (anonymous):

2 over 3 (x − 4) = 2x SOLVE FOR X

OpenStudy (campbell_st):

is it \[\frac{2}{3(x - 4)} = 2x\] or \[\frac{2}{3} (x - 4) = 2x\]

OpenStudy (anonymous):

THE SECOND ONE U WROTE

OpenStudy (anonymous):

R U THERE?

OpenStudy (anonymous):

OK I SEE U R

OpenStudy (anonymous):

AND CAN U SHOW ME HOW U DID IT?

OpenStudy (campbell_st):

ok... so multiply both sides of the equation by 3 to get rid of the 3 in the denominator of the fraction \[3 \times \frac{2}{3} (x - 4) = 3 \times 2x\] so you will have 2(x - 4) = 6x you now need to distribute 2(x - 4)

OpenStudy (campbell_st):

can you do that?

OpenStudy (anonymous):

YA I GOT 2X-8=6X THEN WHAT

OpenStudy (campbell_st):

ok... so subtract 2x from both sides... what do you get..?

OpenStudy (campbell_st):

2x - 8 = 6x -2x -2x --------------

OpenStudy (anonymous):

-8=4X RIGHT?

OpenStudy (anonymous):

THAT IS RIGHT RIGHT?

OpenStudy (campbell_st):

correct... last step divide both sides by 4 to get your answer.